• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name E. Cisbani

    Log entry time 08:01:46 on June20,2007

    Entry number 206029

    keyword=Owl shift summary

    Owl Shift summary

    00:00 Running C12 with radiator in (position 5) at 30 uA, run 1378 at about 500 kevents

    00:38 MCC informs that beam will be off for a couple of minutes; pause the run 1379 failed; then we stop the run.

    00:50 Beam back, new run (1380) started.

    04:15 Found TS0 frozen; reset the right Trigger Supervisor Crate via Hall A General Tools GUI (alexandre advice) and startcoda (see log entry 206020)

    06:40 MCC informs: lost 3 IOC, no beam for a while. Stop run 1385

    07:20 Beam back, run 1386 started

    Left/Q3 drifted few times at about 2.784 GeV/c, each time for about 5 minutes.